OpenStudy (anonymous):

Find the length of the loop for the given curve: x = 12t - 4t^(3) , y = 12t^(2)

myininaya (myininaya):

\[L=\int\limits_{upperintersection}^{lowerintersection}\sqrt{(x')^2+(y')^2} dt\] I believe that is what I remember do you know how to find the intersections?

OpenStudy (anonymous):

no I dont. To be exact my main problem is finding the values of t that I should use for my limits of integration which I guess are just the intersection point.

OpenStudy (anonymous):

points**

OpenStudy (asnaseer):

yes - the loop in question begins and ends at x=0 (see here: http://www.wolframalpha.com/input/?i=x%3D12t-4t%5E3%2C+y%3D12t%5E2) so find the values for t for which x=0 - they will form the lower and upper limits of the integral referred to by @myininaya

OpenStudy (anonymous):

after plugging in all my stuff I get some thing like this: \[\int\limits_{t=0}^{t=\sqrt{3}}\sqrt{144t ^{4}+288t ^{2}+144}dt\]

OpenStudy (anonymous):

I squred the derivatives and collected liked terms inside the radical

OpenStudy (asnaseer):

yup that looks correct - you should be able to spot that 144 can be factored out. that should leave you with a simpler equation that should also be factorable.

OpenStudy (asnaseer):

the square root in the integral will disappear if you spot how to factorize this.

myininaya (myininaya):

sorry i had a thing i had my limits backwards lol

myininaya (myininaya):

ok so did you guys find out how to find the intersections? i can show you how dedesigns

OpenStudy (asnaseer):

also @DEdesigns57 - your limits are not right. should be from minus square root 3 to plus square root 3.

myininaya (myininaya):

well he could double it instead

OpenStudy (asnaseer):

remember:\[x^2=3\implies x=\pm\sqrt{3}\]

OpenStudy (anonymous):

why? it seems the (0,0) is the other intersection point.

OpenStudy (asnaseer):

we are trying to find the start and end limits of 't' - not x.

OpenStudy (asnaseer):

you should have ended up with something like:\[t^2=3\implies t=\pm\sqrt{3}\]

OpenStudy (anonymous):

if t = 0 than (x=0,y=0) though...

OpenStudy (asnaseer):

that only gives you half the curve - so as @myininaya said, you could do 0 to square root 3 and then double the answer.

OpenStudy (anonymous):

and when I look at the graph (0,0) and (0,36) are the intersection points. so I just found the values of t that made that happen. What am I doing wrong?

myininaya (myininaya):

ok if we look at the graph we will see two intersection clearly x is 0 when both intersection happen so x=12t-4t^3 => we have 0=12t-4t^3 at the intersections 0=4t(3-t^2) => t=0, pm sqrt(3) and if we plug these into y we would get y=0 and y=36 And our graph confirms these are the right points of intersection: (0,0) and (0, 36) --------------- based on your graphs symmetry I belive you could just double what you have

myininaya (myininaya):

and i'm late you guys already said this

OpenStudy (asnaseer):

thanks for your assistance @myininaya - always appreciated :-)

OpenStudy (asnaseer):

ok @DEdesigns57 - are you following the explanations above?

OpenStudy (anonymous):

so its not from t=0 to t=sqrt(3)?

OpenStudy (asnaseer):

let me draw it to try and help

OpenStudy (asnaseer):